The measure of two supplementary angles are in the Ratio 11:13. Find the measure of each angle.

Answer by oberobic(2304)

The two supplementary angles total 180 degrees by definition.
They have a ratio of 11:13, so we can define:
11x + 13x = 180
24x = 180
x = 180/24
x = 15/2
.
11x = 11*15/2 = 165/2 = 82.5
13x = 13*15/2 = 195/2 = 97.5
.
Do they total 180 degrees?
82.5 + 97.5 = 180.
Yes.
.
Is their ratio 11:13?
82.5/97.5 = 11/13.
Correct.
.
The supplementary angles are 82.5 and 97.5 degrees.
.
Done.